# Predicting from a Different Perspective in Re-ranking Model for Inductive Knowledge Graph Completion

### v1

@@ Abstract
Rule-induction models have been shown great power in the inductive setting of knowledge graph completion. In this setting, the models are tested on a knowledge graph entirely composed of unseen entities. These models learn relation patterns as rules by utilizing subgraphs. The same input but different rules cause differences in the model’s predictions. In this paper, we focus on this behavior of the model. We propose a re-ranking-based model called ReDistLP (Re-ranking with a Distinct Model for Link Prediction). This model enhances the effectiveness of re-ranking by leveraging the difference in the predictions between the initial retriever and the re-ranker. ReDistLP outperforms the state-of-the-art methods in 2 out of 3 datasets.
@@ 1 Introduction
Knowledge graphs (KGs) store collections of facts in the form of triples. Each triple $(h,r,t)$ represents a relation $r$ between a head entity $h$ and a tail entity $t$. Various KGs have been proposed, including Freebase (Bollacker et al., 2008) for real-world information and GeneOntology (Consortium, 2004) for biological data. These KGs are used in many applications, such as question answering and recommendation systems. Since many KGs are inherently incomplete, it is essential to complete the missing information. Link prediction is a task to predict a missing link in a KG. This completion process improves the KG itself and the usefulness of applications using KGs.
@@
Knowledge Graph Embedding (KGE) such as TransE (Bordes et al., 2013), RotatE (Sun et al., 2019), and ConvE (Dettmers et al., 2018), is the most dominant method for link prediction. Most KGE methods are designed for the transductive setting, where the test entities are a subset of the training entities. These models may not produce meaningful embeddings for unseen entities in the inductive setting. In the inductive setting, entities for the test are not present in the training, while relations are the same (Teru et al., 2020). This setting requires the model to have high inductive ability. Inductive ability is the capability to generalize and make accurate predictions for unseen entities based on the learned patterns and rules from the training data.
@@
GraIL (Teru et al., 2020) and BERTRL (Zha et al., 2022) are designed for the inductive setting. They learn the probabilistic logical rules of relation patterns using subgraphs around an input triple. These models demonstrate a high inductive ability because such rules are entity-independent.
@@
Our empirical findings reveal that BERTRL tends to predict different entities when distinct rules are provided as input. The model trained with the longer rules predicts entities within narrower regions on a KG (Section 4). Motivated by this observation, we propose ReDistLP (Re-ranking with a Distinct Model for Link Prediction), which leverages this phenomenon to improve accuracy. ReDistLP introduces a re-ranking pipeline. The re-ranking pipeline obtains a pool of candidate documents through an inexpensive approach and subsequently re-evaluates them using a more expensive model (Wang et al., 2011). This approach achieves both low computational cost and high accuracy. While existing research has mainly focused on improving the recall of the initial retriever or the performance of the re-ranker, this study focuses on the difference in the predictions between them. We show theoretical insights to improve the accuracy of the re-ranking pipeline. When selecting a model as the re-ranker, choosing one that predicts different entities compared to the initial retriever can improve accuracy. We train three variants of BERTRL, each incorporating rules of different lengths as additional inputs during training and inference. We construct ReDistLP using these variants of BERTRL. ReDistLP achieves the best performance in 2 out of 3 datasets.
The rest of this paper is organized as follows. In Section 2, we describe some related work. In Section 3, we show theoretical insights for maximizing the effect of re-ranking. Section 4 analyzes how the model makes different predictions based on distinct rules. In Section 5, we present our ReDistLP model and provide a detailed explanation of its components. In Section 6, we evaluate the performance of ReDistLP on three benchmark datasets. Section 7 concludes the paper.
@@ 2 Related Works and Backgrounds
2.1 Problem Definition
Formally, a knowledge graph $G$ is defined as a tuple $(\mathcal{E},\mathcal{R},\mathcal{T})$, where $\mathcal{E}$ is a set of entities, $\mathcal{R}$ is a set of relations and $\mathcal{T}$ is a set of triples $\{(h_{i},r_{i},t_{i})\}$. Here, $h_{i},t_{i}\in\mathcal{E}$ denotes the subject and the object, respectively, $r_{i}\in\mathcal{R}$ denotes the relation (Teru et al., 2020).
In the inductive setting of link prediction, we have a training graph $G_{{\rm train}}=(\mathcal{E}_{{\rm train}},\mathcal{R},\mathcal{T}_{{\rm train}})$ and a inference graph $G_{{\rm inf}}=(\mathcal{E}_{{\rm inf}},\mathcal{R},\mathcal{T}_{{\rm inf}})$. The entity sets of $G_{{\rm train}}$ and $G_{{\rm inf}}$ are mutually exclusive ($\mathcal{E}_{train}\cap\mathcal{E}_{inf}=\emptyset$).
Formally, most models designed for the link prediction task assign scores to each entity in response to an input query such as $(h,r)$, and the higher score signifies a greater likelihood of a relevant or accurate link.
@@ 2.2 Knowledge Graph Completion Models
Embedding-based Models
Embedding-based models such as TransE (Bordes et al., 2013), RotatE (Sun et al., 2019), and ConvE (Dettmers et al., 2018), learn the embedding of entities and relations in continuous spaces. These methods are trained to ensure that the operation of the relation on the subject results in the object in the vector space. KG Embedding-based models capture the structural properties of the KG within this embedding space. However, most embedding-based models are transductive; They do not produce appropriate embeddings for unseen entities during the training. Some methods, such as those proposed by (Wang et al., 2019; Hamaguchi et al., 2017), generate embeddings for unseen entities by aggregating the embeddings of neighboring entities learned during training. However, these methods cannot be applied in the inductive setting that handles a KG entirely composed of unseen entities. MorsE (Chen et al., 2022) and StATIK (Markowitz et al., 2022) are embedding-based models that can predict in the inductive setting. MorsE introduces meta-learning to transfer knowledge to unseen entities. StATIK uses a transformer and a message-passing neural network as the encoder, with TransE as the decoder.
Rule-induction Methods and Subgraph-induction Methods
Rule-induction methods learn probabilistic logical rules from a KG. This method inherently can solve the inductive setting since such rules are independent of entities. Here is an example of rules
$(x,{\rm president\_of},y)\land(z,{\rm capital\_of},y)$
$\to(x,{\rm work\_at},z).$ (1)
This rule denotes that if $x$ is the president of $y$, and $z$ is the capital of $y$, then $x$ works at $z$. Once such rules are learned, they can be used to generalize unseen entities. AMIE (Galárraga et al., 2013) and RuleN (Meilicke et al., 2018) mind these rules while reducing the search space. NeuralLP (Yang et al., 2017) and DRUM (Sadeghian et al., 2019) learn such rules in an end-to-end differentiable manner. GraIL (Teru et al., 2020) uses a graph neural network to learn these rules over the subgraphs around the input triple.
Language-based Models
KG-BERT (Yao et al., 2019), GenKGC (Xie et al., 2022), and BERTRL (Zha et al., 2022) are language-based models that use pre-trained language models to utilize the textual information. KG-BERT takes the triple as the textual sequences. It trains BERT to score the plausibility of the input triple. GenKGC takes the missing triple $(h,r,?)$ and generates a missing entity. While KG-BERT requires scoring all possible triples in inference, GenKGC can infer entities directly through generation. Thus, GenKGC offers faster inference times than KG-BERT. However, both KG-BERT and GenKGC have limited inductive ability since they do not utilize subgraph information. Their inductive capacity is solely attributed to the pre-trained language model. In contrast, BERTRL demonstrates a high inductive ability. Similar to KG-BERT, BERTRL is a model designed to identify triples. This model takes the paths connecting head and tail entities, along with the input triple. Its high inductive capability arises from both a pre-trained language model and a subgraph reasoning mechanism. Hence, BERTRL can also be classified as a subgraph-induction method. The formulation for BERTRL is as follows
@@
2.3 Re-ranking Models
The re-ranking pipeline initially retrieves a collection of candidate documents with a less costly model and subsequently re-scores them using a more expensive model (Wang et al., 2011). Usually, the re-ranker model has higher accuracy but entails higher inference costs. This trade-off is crucial for the efficiency and effectiveness of the model. A higher number of candidate documents need to be retrieved to increase the efficacy, but this leads to a decrease in efficiency. Reducing the size of the candidate set can enhance efficiency, as it enables the re-ranker to minimize the time spent on re-ranking the documents; however, this results in a reduction in effectiveness (Chen et al., 2017; MacAvaney et al., 2022).
Various re-ranking models have been proposed to enhance either efficiency or effectiveness. GAR (MacAvaney et al., 2022) dynamically updates the initial pool to improve the recall of the initial retriever. This model repeatedly adds a document similar to the initial pool’s top-scoring ones. The monoBERT (Nogueira et al., 2019) and duoBERT (Nogueira et al., 2019) are BERT-based textual ranking models. The monoBERT scores documents based on binary classification. The duoBERT guesses which sentence is more relevant to the query. (Chen et al., 2017) proposed a cost-aware re-ranking model by introducing the feature costs. This model optimizes latency by selecting feature sets based on the importance and costs of features.
3 Maximize the Effectiveness of Re-ranking
We provide theoretical insights to maximize the effectiveness of re-ranking. Let $score_{m}$ represent the score function in model $m$, and $\bar{t}\in\mathcal{E}$ denote the correct answer. Now we consider an entity set where the score of each entity is greater than or equal to the score of $\bar{t}$:
$S_{m}(h,r)=\{t\in\mathcal{E}|score_{m}(h,r,t)\geq score_{m}(h,r,\bar{t})\}.$ (3)
The size of $S_{m}(h,r)$ is equivalent to the performance of model $m$ for query $(h,r)$ because $1/|S_{m}(h,r)|$ corresponds to the Reciprocal Rank (RR). Mean Reciprocal Rank (MRR) is widely used as an evaluation metric in research fields such as information retrieval systems and recommender systems.
Suppose that model $m_{1}$ is a initial retriever, and model $m_{2}$ is a re-ranker. Let us assume the accuracy of $m_{2}$ exceeds that of $m_{1}$ ( $|S_{m_{1}}|>|S_{m_{2}}|$). After re-ranking, the predicted entities are given by
$S^{\prime}_{m_{1},m_{2}}(h,r)=\{t\in S_{m_{1}}(h,r)|score_{m_{2}}(h,r,t)\geq score_{m_{2}}(h,r,\bar{t})\}.$ (4)
Given the above, we can infer
$|S^{\prime}_{m_{1},m_{2}}(h,r)|\leq|S_{m_{2}}(h,r)|<|S_{m_{1}}(h,r)|.$ (5)
This equation indicates that the performance after re-ranking exceeds that of the initial retriever and the re-ranker. Equation (5) can be resolved into the following equation and inequality, depending on the relationship between $S_{m_{1}}$ and $S_{m_{2}}$:
$S_{m_{1}}\subset S_{m_{2}}\implies|S^{\prime}_{m_{1},m_{2}}|=|S_{m_{2}}|.$ (6)
This equation suggests that when $S_{m_{1}}$ is a subset of $S_{m_{2}}$, there will be no improvement in accuracy after re-ranking:
$S_{m_{1}}\not\subset S_{m_{2}}\implies|S^{\prime}_{m_{1},m_{2}}|<|S_{m_{2}}|.$ (7)
This equation implies that when $S_{m_{1}}$ is not a subset of $S_{m_{2}}$, the accuracy after re-ranking will be improved compared to that of the re-ranker.
Furthermore, by selecting models based on the next equation, re-ranking effectiveness is maximized through the minimization of $S_{m_{1},m_{2}}$:
$\operatorname*{argmax}_{\;\;m_{1},m_{2}}\;|S_{m_{1}}\setminus S_{m_{2}}|.$ (8)
From the above observations, it is ineffective to select the models that make almost the same predictions with each other; it is crucial to choose the models that offer distinct predictions from one another.
4 Analyzing Model Behavior Under Different Rule Sets
In this section, we demonstrate that BERTRL predicts distinct entities when provided with different rules as input. We train BERTRL with 3-hop, 2-hop, and 1-hop rules, respectively, and analyze the differences in their prediction sets.
To investigate the differences in predictions between the variant models directly, we summarize the intersection ratio of top-$n$ prediction sets in Table 1. The results show that the prediction sets differ significantly depending on the number of hops used in the training rules. The predictions of BERTRL trained with 1-hop and 2-hop rules are the most similar, while those trained with 1-hop and 3-hop rules are the least similar.
@@ 4 Analyzing Model Behavior Under Different Rule Sets
Table 2: $n$-hop Reachability Between Entities in Top-10 Predicted Sets
@@
$n=2$
BERTRL 3-hop 0.338 0.639 0.894
BERTRL 2-hop 0.178 0.612 0.899
BERTRL 1-hop 0.0624 0.283 0.538
$n=3$
BERTRL 3-hop 0.441 0.860 0.980
BERTRL 2-hop 0.239 0.777 0.958
BERTRL 1-hop 0.110 0.487 0.740
$n=5$
BERTRL 3-hop 0.577 0.958 0.996
BERTRL 2-hop 0.392 0.910 0.992
BERTRL 1-hop 0.282 0.797 0.959
Furthermore, we observe that the longer the number of hops, the more the predicted entities tend to be concentrated locally on the graph. Figure 1 shows the top-10 predicted entities and the paths that connect them within 3-hop. Table 2 shows $n$-hop reachability between entities in top-10 predicted sets.
(a) Hop 1 (WN18RR)
(b) Hop 1 (FB15k-237)
(c) Hop 1 (NELL-995)
(d) Hop 2 (WN18RR)
(e) Hop 2 (FB15k-237)
(f) Hop 1 (NELL-995)
(g) Hop 3 (WN18RR)
(h) Hop 2 (FB15k-237)
(i) Hop 3 (NELL-995)
@@
From Table 2 and Figure 1, it can be observed that: (1) Reachability in predicted entities increases with the number of hops of rules. (2) BERTRL with a higher number of hops predicts entities within narrower regions on a KG, while predicted entities by BERTRL with fewer hops spread more widely on a KG.

### v2

@@ Abstract
Rule-induction models have demonstrated great power in the inductive setting of knowledge graph completion. In this setting, the models are tested on a knowledge graph entirely composed of unseen entities. These models learn relation patterns as rules by utilizing subgraphs. Providing the same inputs with different rules leads to differences in the model’s predictions. In this paper, we focus on the behavior of such models. We propose a re-ranking-based model called ReDistLP (Re-ranking with a Distinct Model for Link Prediction). This model enhances the effectiveness of re-ranking by leveraging the difference in the predictions between the initial retriever and the re-ranker. ReDistLP outperforms the state-of-the-art methods in 2 out of 3 benchmarks.
Keywords Knowledge Graph  $\cdot$ Inductive Knowledge Graph Completion  $\cdot$ Re-ranking
@@ 1 Introduction
Knowledge graphs (KGs) store collections of facts in the form of triples. Each triple $(h,r,t)$ represents a relation $r$ between a head entity $h$ and a tail entity $t$. Various KGs have been proposed, including Freebase (Bollacker et al., ) for real-world data and GeneOntology (Consortium, 2004) for biological data. These KGs are used in many applications, such as question-answering and recommendation systems; however, since many KGs are inherently incomplete, completing the missing information is essential. One way to do this is by link prediction, which is a task that predicts missing links in a KG. Such Knowledge Graph Completion (KGC) improves the KG itself and enhances the usefulness of applications using the KG
@@
Knowledge Graph Embedding (KGE), such as TransE (Bordes et al., ) and RotatE (Sun et al., ), is the most dominant method for link prediction. Most KGE methods are designed for the transductive setting, where the test entities are a subset of the training entities. These models may not produce meaningful embeddings for unseen entities in the inductive setting. In the inductive setting, where entities for testing are absent in the training data, while the relations remain the same(Teru et al., ). This setting requires the model to have a high inductive ability. Inductive ability is the capability to generalize and make accurate predictions for unseen entities based on the learned patterns and rules from the training data.
@@
GraIL (Teru et al., ) and BERTRL (Zha et al., ) are designed for the inductive setting. They learn the probabilistic logical rules of relation patterns using subgraphs around an input triple. These models demonstrate high inductive ability because such rules are entity-independent.
@@
Our empirical findings reveal that BERTRL tends to predict different entities when distinct rules are provided as input. The model trained with the longer rules predicts entities within narrower regions on a KG.Motivated by this observation, we propose ReDistLP (Re-ranking with a Distinct Model for Link Prediction), which leverages this phenomenon to improve accuracy. ReDistLP introduces a re-ranking pipeline using two variants of BERTRL that are trained with different lengths of rules. The re-ranking pipeline obtains a pool of candidate documents through an inexpensive approach and subsequently re-evaluates them using a more expensive model (Wang et al., ). This approach achieves both low computational cost and high accuracy. While existing research has mainly focused on improving the recall of the initial retriever or the performance of the re-ranker, this study focuses on the difference in the predictions between them. We show theoretical insights to improve the accuracy of the re-ranking pipeline. When selecting a model as the re-ranker, choosing one that predicts different entities compared to the initial retriever can improve accuracy. We train three variants of BERTRL, each incorporating rules of different lengths as additional inputs during training and inference. We construct ReDistLP using these variants of BERTRL. ReDistLP achieves the best performance in 2 out of 3 benchmarks.
@@ 2.1 Knowledge Graph Completion Models
2.1.1 Rule-induction Methods.
Rule-induction methods learn probabilistic logical rules from a KG. This method inherently can solve the inductive setting since such rules are independent of entities. Here is an example of a rule
$(x,{\rm president\_of},y)\land(z,{\rm capital\_of},y)\to(x,{\rm work\_at},z).$ (1)
This rule denotes that if $x$ is the president of $y$, and $z$ is the capital of $y$, then $x$ works at $z$. Once such rules are learned, they can be used to generalize unseen entities. AMIE (Galárraga et al., ) and RuleN (Meilicke et al., ) mine these rules while reducing the search space. NeuralLP (Yang et al., ) and DRUM (Sadeghian et al., ) learn such rules in an end-to-end differentiable manner. GraIL (Teru et al., ) uses a graph neural network to learn these rules over the subgraphs around the input triple.
2.1.2 Language-based Models.
KG-BERT (Yao et al., 2019) and BERTRL (Zha et al., ) are language-based models that use pre-trained language models to utilize the textual information. KG-BERT takes the triple as textual sequences. It trains BERT to score the plausibility of the input triple. KG-BERT has limited inductive ability since they do not utilize subgraph information. Their inductive capacity is solely attributed to the pre-trained language model. In contrast, BERTRL demonstrates high inductive ability. Similar to KG-BERT, BERTRL is a model designed to identify triples. This model takes the paths connecting head and tail entities and the input triple. Its high inductive capability arises from both a pre-trained language model and a subgraph reasoning mechanism. Hence, BERTRL can also be classified as a subgraph-induction method. The formulation for BERTRL is as follows
@@
3 Analysis of Model Behavior Under Different Rule Sets
This section demonstrates that BERTRL predicts distinct entities when different rules are provided as input. We train BERTRL with 3-hop, 2-hop, and 1-hop rules, respectively, and analyze the differences in their prediction sets.
We investigate the dissimilarities in predictions between variant models from the intersection ratio of top-$n$ prediction sets in Table 1. The results show that the prediction sets significantly depend on the number of hops used in the training rules. The predictions of BERTRL trained with 1-hop and 2-hop rules are the most similar, while those trained with 1-hop and 3-hop rules are the least similar.
@@ 3 Analysis of Model Behavior Under Different Rule Sets
Table 2: $k$-hop reachability between entities in top-10 predicted sets
@@
$k=2$
3-hop 0.338 0.639 0.894
2-hop 0.178 0.612 0.899
1-hop 0.0624 0.283 0.538
$k=3$
3-hop 0.441 0.860 0.980
2-hop 0.239 0.777 0.958
1-hop 0.110 0.487 0.740
$k=5$
3-hop 0.577 0.958 0.996
2-hop 0.392 0.910 0.992
1-hop 0.282 0.797 0.959
Furthermore, we observe that the longer the number of hops, the more the predicted entities tend to be concentrated locally on the KG. Figure 1 shows the top-10 predicted entities and the paths that connect them within 3-hop. Table 2 shows $n$-hop reachability between entities in top-10 predicted sets.
In Table 2 and Figure 1, it can be observed that: (1) Reachability in predicted entities increases with the number of hops of rules. (2) BERTRL with a higher number of hops predicts entities within narrower regions on a KG, while predicted entities by BERTRL with fewer hops spread more widely on a KG.
(a) 1-Hop (FB15k-237)
(b) 2-Hop (FB15k-237)
(c) 3-Hop (FB15k-237)
@@ 4 Proposed Approach
This section first provides insight into enhancing re-ranking effectiveness for our model. Then, we introduce our model of ReDistLP and explain how to re-rank the candidate entities for our model.
4.1 Maximize the Effectiveness of Re-ranking
We provide theoretical insights to maximize re-ranking effectiveness through fuzzy sets. The predictions of a model $m$ can be noted using fuzzy sets.
$S_{m}=\{(t,\mu_{S_{m}}(t))|t\in\mathcal{E}\},$ (3)
where $\mu_{S_{m}}$ is a membership function that assigns to item membership degree between 0 and 1 (Zimmermann, 2010). We use the model predictions as the membership function, i.e., $\mu_{S_{m}}(t)=score_{m}(h,r,t)$.
Let $\bar{t}\in\mathcal{E}$ denote the correct answer for the input triple $(h,r,?)$. Now, we consider $\alpha$-cut for the fuzzy set. The $\alpha$-cut of a fuzzy set is a classical set containing all the elements with a membership degree greater than or equal to a value $\alpha$ (Zimmermann, 2010).
$S_{m}^{\geq\mu_{S_{m}}(\bar{t})}=\{t\in\mathcal{E}|\mu_{S_{m}}(t)\geq\mu_{S_{m}}(\bar{t})\}.$ (4)
The size of $S_{m}^{\geq\mu_{m}(\bar{t})}$ is equivalent to the performance of model $m$ for query $(h,r,?)$, because $1/|S_{m}^{\geq\mu_{m}(\bar{t})}|$ corresponds to the Reciprocal Rank (RR). The Mean Reciprocal Rank (MRR) is widely used as an evaluation metric in research fields such as information retrieval systems, recommender systems, and KGC.
Suppose that model $m_{1}$ is a initial retriever, and model $m_{2}$ is a re-ranker. Let us assume the accuracy of $m_{2}$ exceeds that of $m_{1}$ ( $|S^{\geq\mu_{S_{m_{1}}}(\bar{t})}_{m_{1}}|>|S^{\geq\mu_{S_{m_{2}}}(\bar{t})}_{m_{2}}|$). After re-ranking, the set of predicted entities is given by
$S^{\prime}_{m_{1},m_{2}}=\{t\in S_{m_{1}}^{\geq\mu_{S_{m_{1}}}(\bar{t})}|\mu_{S_{m_{2}}}(t)\geq\mu_{S_{m_{2}}}(\bar{t})\}.$ (5)
Given the above, we can infer
$\bigl{|}S^{\prime}_{m_{1},m_{2}}\bigr{|}\leq\bigl{|}S_{m_{2}}^{\geq\mu_{S_{m_{2}}}(\bar{t})}\bigr{|}<\bigl{|}S_{S_{m_{1}}}^{\geq\mu_{S_{m_{1}}}(\bar{t})}\bigr{|}.$ (6)
This equation indicates that the performance after re-ranking exceeds that of the initial retriever and the re-ranker. Equation (6) can be resolved into the following equation and inequality, depending on the relationship between $S_{m_{1}}^{\geq\mu_{S_{m_{1}}}(\bar{t})}$ and $S_{m_{2}}^{\geq\mu_{S_{m_{2}}}(\bar{t})}$:
$S_{m_{1}}^{\geq\mu_{S_{m_{1}}}(\bar{t})}\subset S_{m_{2}}^{\geq\mu_{S_{m_{2}}}(\bar{t})}\implies\bigl{|}S^{\prime}_{m_{1},m_{2}}\bigr{|}=\bigl{|}S_{m_{2}}^{\geq\mu_{S_{m_{2}}}(\bar{t})}\bigr{|}$ (7)
This equation suggests that when $S^{\geq\mu_{S_{m_{1}}}(\bar{t})}_{m_{1}}$ is a subset of $S^{\geq\mu_{S_{m_{2}}}(\bar{t})}_{m_{2}}$, there will be no improvement in accuracy after re-ranking. On the other hand, if $S_{m_{1}}^{\geq\mu_{S_{m_{1}}}(\bar{t})}$ is not a subset of $S_{m_{2}}^{\geq\mu_{S_{m_{2}}}(\bar{t})}$, the following equation holds, implying that the accuracy after re-ranking will be higher than that of the re-ranker alone:
$S_{m_{1}}^{\geq\mu_{S_{m_{1}}}(\bar{t})}\not\subset S_{m_{2}}^{\geq\mu_{S_{m_{2}}}(\bar{t})}\implies\bigl{|}S^{\prime}_{m_{1},m_{2}}\bigr{|}<\bigl{|}S_{m_{2}}^{\geq\mu_{S_{m_{2}}}(\bar{t})}\bigr{|}$ (8)
Furthermore, by selecting models based on the next equation, re-ranking effectiveness is maximized through the minimization of $S^{\prime}_{m_{1},m_{2}}$:
$\operatorname*{argmax}_{\;\;m_{1},m_{2}}\;\Bigl{|}S_{m_{1}}^{\geq\mu_{S_{m_{1}}}(\bar{t})}\setminus S_{m_{2}}^{\geq\mu_{S_{m_{2}}}(\bar{t})}\Bigr{|}$ (9)
From the above observations, it is ineffective to select the models that make almost the same predictions; it is crucial to choose the models that offer distinct predictions.
4.2 Model Details